OpenStudy (anonymous):

solve the equation in the specified interval cosx=-0.5 360<_x< 720

OpenStudy (anonymous):

Our referance angle is 60 degrees. Thus the general equation to x is x = 60+360n *or* x=-60+360n. Trying different n we get 420 and 680